# Electronic correlation effects in the response of graphene and MoS2 monolayers to the impact of highly-charged ions

### v1

@@ 1 Introduction
The dynamics of correlated fermions driven out of equilibrium by external forces, such as laser excitation or potential quenches, have attracted high interest in recent years. Examples are the diffusion of fermionic atoms following a confinement quench, e.g. [1, 2], and laser excitation of correlated materials in condensed matter [3]. A third example is the response of two-dimensional (2D) quantum materials to the impact of highly-charged ions [4, 5, 6]. This process leads to a very strong excitation of the solid which is very localized in space (on a few lattice sites) and time (a few femtoseconds). Studying the response of the material to the excitation and the thermalization provides fundamental insights into the electronic and optical properties of the material and may, ultimately, lead to novel technologies in nanoelectronics, such as petahertz switching [7].
@@
A key approximation is Hartree-Fock (HF) which includes many-particle effects via an effective potential (mean field) created by all particles [8, 9, 10]. HF is computationally advantageous and can, in some cases, already capture many of the interaction-based effects in many-body systems. However, correlation effects are neglected which are important, e.g. for two-electron collisions, Auger ionization or interatomic Coulombic decay (ICD) [11]. The most popular approach to include correlations is time-dependent density functional theory (TDDFT) [12]. In this method an exchange-correlation functional of the particle density is introduced in order to include deviations from the Hartree solution. While TDDFT allows to efficiently treat real materials, its capability to describe quasiparticle excitations, correlation effects, and their dynamics is contingent to a suitable choice of the exchange-correlation kernel, for which systematic approximation schemes remain elusive.
@@
Successful alternative methods are reduced density operators [13, 14, 15] and non-equilibrium Green function theory (NEGF). Here we focus on NEGF because it provides systematic approximation schemes, via Feynman diagrams, to correlations and, moreover, direct access to dynamical and spectral properties of the system [16, 17, 13]. However, these advantages come with a high computational effort: cubic scaling of the computation time with the number of time steps, $N_{t}$, as a result of the two time arguments of the one-particle Green function and the memory time integral, limiting the method to short simulations and/or small systems.
@@
These bottlenecks could be overcome to a large extent with the use of the generalized Kadanoff–Baym ansatz (GKBA) [18, 19, 13, 20] where the time propagation of the Green functions is restricted to the time diagonal. In particular, the derivation of the time-local G1–G2 scheme [21, 22, 23] has led to recent breakthroughs in the computational capabilities, for a recent overview, see Ref. [24]. With this scheme it is now possible to perform quantum many-body simulations of non-uniform systems with sub-femtosecond time resolution with time-linear scaling ($\propto N_{t}$) [21]. While the G1–G2 scheme is closely linked to reduced density matrix (RDM) theory derived from the BBGKY-hierarchy [25, 26, 23], it provides direct access to important common NEGF approximation schemes, including the T-matrix or GW-approximations and, therefore, a sound many-body foundation.
@@
While, with the G1–G2 scheme, the CPU time problem of the propagation of correlated electron systems has been overcome, the remaining main computational bottleneck is the costly storage of the two-particle Green function, $G_{2}$, for a recent detailed analysis, see Ref. [24]. The main strategy here is either to reduce the basis dimension or to use a basis where the pair interaction is highly diagonal. An example where this is the case is the Hubbard model or its various extensions, such as the Pariser-Parr-Popel (PPP) model. In recent years, Hubbard-type models were successfully applied to many interesting situations of correlated fermion physics, such as confinement quench-excited diffusion of correlated fermionic atoms [2, 27], laser excitation of graphene [24, 28], and ion stopping in correlated materials [29, 30, 6]. With a proper selection of the model parameters, usually derived from DFT, high accuracy for the electron dynamics can be achieved [2, 27]. Therefore, in the present paper we will use a Hubbard model to describe the interaction of monolayers of quantum materials with a highly-charged ion. A second complementary approach to reduce the size of the interaction matrix and of $G_{2}$ is to apply an embedding approach [17, 31]. The idea is to split the system into two coupled subsystems of which only one (the “main” system that is of primary interest) is treated on an advanced many-body level whereas for the rest (the “environment”) a simpler (and less costly) approximation is being used. In fact, an embedding concept was successfully used to model the neutralization of highly-charged ions [32]. Recently the embedding ansatz has been combined with the G1–G2 scheme such that the time-linear scaling of the full simulation is retained [33, 24].
@@
Thus, in our present analysis we use the G1–G2 scheme, in combination with a time-linear embedding scheme, in order to verify and extend the theoretical results of Ref. [6]. There it was found that graphene and MoS2 show drastic differences in their reaction to the impact of slow highly-charged ions. The most striking one is a 7-fold higher number of secondary electrons that are emitted from a graphene monolayer in comparison to MoS2. The theoretical analysis of Ref. [6] that used NEGF simulations on the Hartree-Fock level explained this strong difference in the behavior of the two materials by the much higher electron mobility in graphene. The consequence was that, in graphene, electron loss due to resonant charge transfer to the approaching was compensated more quickly than in MoS2. However, the reliability of this conclusion remains to be tested.
@@
Therefore, the main goal of the present paper is to study the influence of electronic correlations that were neglected in Ref. [6], on the ion-monolayer interaction [34]. While correlation effects are known to be moderate in graphene, which has no bandgap, in monolayers of semiconducting MoS2, which have a bandgap of around 1.9 eV [35], a significant influence is expected. With this property MoS2 can be regarded as a prototype for many 2D quantum materials with high prospects to applications in nanotechnology, e.g., [36]. So exploring their response to a strong excitation with sub-femtosecond time resolution is of high relevance for other excitation schemes of quantum materials.
@@
This paper is structured as follows. In Sec. 2, we derive and discuss the theoretical framework of the G1–G2 scheme, including the time-local embedding scheme. In Sec. 3 we apply the derived equations to ion stopping in graphene and MoS2 monolayers using Hubbard-type lattice models in combination with a model for the resonant charge transfer between the ion and the monolayer. In Sec. 4 we present the numerical results which include the charge density dynamics in the monolayer during the ion impact. We then compute the electrostatic potential that is produced by this strongly non-uniform charge density in the the monolayer which has been found to be crucial for the emission of secondary electrons [6]. Finally, we present results for the spectral properties during the ion impact. The results are summarized in Sec. 5 where we also give an outlook for future developments.
@@ 2.2 Non-equilibrium Green functions
In this section, we give a brief introduction into NEGFs and outline the derivation of the G1–G2 scheme. More information and details on the formalism can be found in Refs. [22, 23, 24].
@@
In this expression $h^{\mathrm{0}}(t)$ denotes the one-particle Hamiltonian, containing the kinetic energy as well as external potentials, such as the Coulomb potential of the approaching, highly-charged ion [29, 30]. $w_{ijkl}(t)$ are the matrix elements of the pair interaction potential between electrons. The time dependence reflects the adiabatic switch-on of the potential what allows us to start the simulation from a non-interacting initial state which is known [16]. The definition of the one-particle Green function is given by
@@
The equations of motion for $G$ on the contour are the Keldysh-Kadanoff-Baym equations (KBEs) [37, 39]
@@
the equation of motion of which couples to the three-particle Green function (Martin-Schwinger-hierarchy [40]). The other way [Eq. (2)] eliminates the two-particle quantity in favor of the one-particle self-energy $\Sigma$. For practical applications we switch to real times and consider the correlation functions
@@
In addition, we have access to spectral properties which require the Fourier transform, with respect to the difference time $\tau=t^{\prime}-t$, of the difference, $G^{\mathrm{>}}_{ij}(t,t^{\prime})-G^{\mathrm{<}}_{ij}(t,t^{\prime})$ [17, 13]. However, there is also a cost to this advantage, namely the computationally expensive contour-time integral on the r.h.s. of the KBEs. The computation of this integral results in a cubic scaling of the CPU time with the propagation duration, making long simulations impossible, except for very small systems.
@@ 2.3 HF-GKBA and G1–G2 scheme
$i\hbar\frac{\partial}{\partial t}G_{ij}^{\mathrm{<}}(t)-\big{[}\bm{h}^{\mathrm{HF}},\bm{G}^{\mathrm{<}}\big{]}_{ij}(t)=i\hbar\sum_{k}\int_{t_{0}}^{\mathrm{t}}\mathrm{d}\,\bar{t}\,\left[\Sigma^{\mathrm{SOA,>}}_{ik}(t,\bar{t}\,)G_{kj}^{\mathrm{<}}(\bar{t},t)-\Sigma^{\mathrm{SOA,<}}_{ik}(t,\bar{t}\,)G_{jj}^{\mathrm{>}}(\bar{t},t)\right]\,$,
where we denoted $A,B]_{ij}:=\sum_{k}(A_{ik}B_{kj}-B_{ik}A_{kj})$. Since the evolution on the time diagonal still depends on the off-diagonal elements, under the time integral, the equation of motion (EOM) is not fully time-diagonal and the costly memory integral still occurs.
For this reason, a common way to approximate the full two-time solution is to use a reconstruction of the off-diagonal part of the one-particle Green function [18]
@@
The reconstruction already represents an approximation since additional terms, involving time integrals as well as the self-energy, are neglected [18]. The ansatz (5) for the off-diagonal parts is commonly known as the Generalized Kadanoff–Baym Ansatz (GKBA). To apply this reconstruction, we also need to know the time-off-diagonal elements of the propagators $G^{\mathrm{R/A}}(t,t^{\prime})$. The most common choices are either the free GKBA or the HF-GKBA, where the propagators are the free and the Hartree-Fock propagators, respectively. In SOA, this approximation reduces the computational cost significantly (to quadratic in the number of time steps), but for higher-order self-energies the cubic scaling remains. Furthermore, due to the use of HF propagators, spectral properties do not contain correlation effects.
As described in Ref. [21] and further expanded upon in Refs. [24, 22, 23], one can, in fact, by using the GKBA, completely remove the memory integral from the EOM for the Green function on the time diagonal. This is achieved by utilizing the KBEs with collision integrals defined in terms of the two-particle Green function, Eq. (1), instead of the self-energy formulation of Eq. (2). The strategy is fairly straightforward:
@@
The result is the G1–G2 scheme which has been proven to exhibit linear scaling in the propagation time at the cost of propagating, in addition to $G^{<}$, the two-particle Green function [21].
@@
where, in Eq. (8), we introduced the two-particle Hartree and Fock Green functions. Finally, by applying the HF-GKBA and differentiating the expression of $G^{\mathrm{(2)}}$ with respect to time we obtain the coupled set of EOMs for the G1–G2 scheme in SOA [28]
@@
Using the same strategy, one can find EOMs for all higher-order self-energies, as demonstrated in [23]. There, comparisons between the RDM and the G1–G2 approach were performed. The authors conclude that both methods are equivalent in HF, SOA and TPP approximation with and without exchange, meaning that for all known NEGF self-energies one can find an equivalent G1–G2 formulation. Since approximations stemming from self-energies can be better understood in terms of the diagrams they include, it is very useful to be able to assign them to the time-diagonal contributions. At the same time, the G1–G2 scheme allows for additional many-body approximations for which no direct single-particle self-energies are known. An example is the dynamically screened ladder (DSL) approximation, for details, see Ref. [23].
In this work we will focus on the HF and SOA self-energies and make use of the G1–G2 framework which has been tested multiple times for simple Hubbard chains and more complex finite graphene clusters [22, 23].
@@ 2.4 NEGF embedding approach
In the following, we focus on a special case of the G1–G2 scheme, namely the description of a total system consisting of two distinct parts which will be called “system” and “environment”. Possible applications for such a framework are the description of electron dynamics in a crystal with metal contacts or, as we will discuss later in Sec. 3, an ion colliding with a graphene or MoS2 monolayer. Of course the G1–G2 scheme could be used to describe the full system in detail but that might require unnecessary computation cost since the system and the environment might evolve on completely different time scales and correlation effects might not be of equal importance in the two parts. The goal of the embedding scheme is, therefore, to be able to employ different approximations for the correlations in both, the system and the environment. For example, in the ion stopping scenario, one might want to use a higher-order self-energy, such as SOA, in the system (the target), while treating the ion without correlations altogether. The derivation of such a framework from the NEGF perspective is demonstrated, e.g., in Refs. [17, 32, 33].

### v2

@@ 1 Introduction
The dynamics of correlated fermions driven out of equilibrium by external forces, such as laser excitation or potential quenches, have attracted high interest in recent years. Examples are the diffusion of fermionic atoms following a confinement quench, e.g. [schneider_np_12, schluenzen_prb16], and laser excitation of correlated materials in condensed matter [joost_pss_18]. A third example is the response of two-dimensional (2D) quantum materials to the impact of highly-charged ions [gruber_ultrafast_2016, Niggas_phys-com_21, niggas_prl_22]. This process leads to a very strong excitation of the solid which is very localized in space (on a few lattice sites) and time (a few femtoseconds). Studying the response of the material to the excitation and the thermalization provides fundamental insights into the electronic and optical properties of the material and may, ultimately, lead to novel technologies in nanoelectronics, such as petahertz switching [hommelhoff_nat_22].
@@
A key approximation is Hartree-Fock (HF) which includes many-particle effects via an effective potential (mean field) created by all particles [hartree_wave_1928, fock_naherungsmethode_1930, slater_simplified_1954]. HF is computationally advantageous and can, in some cases, already capture many of the interaction-based effects in many-body systems. However, correlation effects are neglected which are important, e.g. for two-electron collisions, Auger ionization or interatomic Coulombic decay (ICD) [averbukh_prl_04]. The most popular approach to include correlations is time-dependent density functional theory (TDDFT) [runge_density-functional_1984]. In this method an exchange-correlation functional of the particle density is introduced in order to include deviations from the Hartree solution. While TDDFT allows to efficiently treat real materials, its capability to describe quasiparticle excitations, correlation effects, and their dynamics is contingent to a suitable choice of the exchange-correlation kernel, for which systematic approximation schemes remain elusive.
@@
Successful alternative methods are reduced density operators [bonitz_qkt, kremp-etal.97ap, donsa_prr_23] and non-equilibrium Green function theory (NEGF). Here we focus on NEGF because it provides systematic approximation schemes, via Feynman diagrams, to correlations and, moreover, direct access to dynamical and spectral properties of the system [schluenzen_jpcm_19, stefanucci_many-body_2013, bonitz_qkt]. However, these advantages come with a high computational effort: cubic scaling of the computation time with the number of time steps, $N_{t}$, as a result of the two time arguments of the one-particle Green function and the memory time integral, limiting the method to short simulations and/or small systems.
@@
These bottlenecks could be overcome to a large extent with the use of the generalized Kadanoff–Baym ansatz (GKBA) [lipavsky_generalized_1986, bonitz-etal.96jpcm, bonitz_qkt, hermanns_prb14] where the time propagation of the Green functions is restricted to the time diagonal. In particular, the derivation of the time-local G1–G2 scheme [schluenzen_prl_20, joost_prb_20, joost_prb_22] has led to recent breakthroughs in the computational capabilities, for a recent overview, see Ref. [bonitz_pssb23]. With this scheme it is now possible to perform quantum many-body simulations of non-uniform systems with sub-femtosecond time resolution with time-linear scaling ($\propto N_{t}$) [schluenzen_prl_20]. While the G1–G2 scheme is closely linked to reduced density matrix (RDM) theory derived from the BBGKY-hierarchy [bogolyubov_problems_1960, boer_studies_1962, joost_prb_22], it provides direct access to important common NEGF approximation schemes, including the T-matrix or GW-approximations and, therefore, a sound many-body foundation.
@@
While, with the G1–G2 scheme, the CPU time problem of the propagation of correlated electron systems has been overcome, the remaining main computational bottleneck is the costly storage of the two-particle Green function, $G_{2}$, for a recent detailed analysis, see Ref. [bonitz_pssb23]. The main strategy here is either to reduce the basis dimension or to use a basis where the pair interaction is highly diagonal. An example where this is the case is the Hubbard model or its various extensions, such as the Pariser-Parr-Popel (PPP) model. In recent years, Hubbard-type models were successfully applied to many interesting situations of correlated fermion physics, such as confinement quench-excited diffusion of correlated fermionic atoms [schluenzen_prb16, schluenzen_prb17], laser excitation of graphene [bonitz_pssb23, joost_phd_2022], and ion stopping in correlated materials [balzer_prb16, balzer_prl_18, niggas_prl_22]. With a proper selection of the model parameters, usually derived from DFT, high accuracy for the electron dynamics can be achieved [schluenzen_prb16, schluenzen_prb17]. Therefore, in the present paper we will use a Hubbard model to describe the interaction of monolayers of quantum materials with a highly-charged ion. A second complementary approach to reduce the size of the interaction matrix and of $G_{2}$ is to apply an embedding approach [stefanucci_many-body_2013, Bonitz_fcse_19]. The idea is to split the system into two coupled subsystems of which only one (the “main” system that is of primary interest) is treated on an advanced many-body level whereas for the rest (the “environment”) a simpler (and less costly) approximation is being used. In fact, an embedding concept was successfully used to model the neutralization of highly-charged ions [balzer_cpp_21]. Recently the embedding ansatz has been combined with the G1–G2 scheme such that the time-linear scaling of the full simulation is retained [kwok_nl_19, balzer_prb_23, bonitz_pssb23, tuovinen_prl_23] which is of high relevance also for transport simulations.
@@
Thus, in our present analysis we use the G1–G2 scheme, in combination with a time-linear embedding scheme, in order to verify and extend the theoretical results of Ref. [niggas_prl_22]. There it was found that graphene and MoS2 show drastic differences in their reaction to the impact of slow highly-charged ions. The most striking one is a 7-fold higher number of secondary electrons that are emitted from a graphene monolayer in comparison to MoS2. The theoretical analysis of Ref. [niggas_prl_22] that used NEGF simulations on the Hartree-Fock level explained this strong difference in the behavior of the two materials by the much higher electron mobility in graphene. The consequence was that, in graphene, electron loss due to resonant charge transfer to the approaching was compensated more quickly than in MoS2. However, the reliability of this conclusion remains to be tested.
@@
Therefore, the main goal of the present paper is to study the influence of electronic correlations that were neglected in Ref. [niggas_prl_22], on the ion-monolayer interaction [lovato_msc_2024]. While correlation effects are known to be moderate in graphene, which has no bandgap, in monolayers of semiconducting MoS2, which have a bandgap of around 1.9 eV [splendiani_emerging_2010], a significant influence is expected. With this property MoS2 can be regarded as a prototype for many 2D quantum materials with high prospects to applications in nanotechnology, e.g., [wang_electronics_2012]. So exploring their response to a strong excitation with sub-femtosecond time resolution is of high relevance for other excitation schemes of quantum materials.
@@
This paper is structured as follows. In Sec. 2, we derive and discuss the theoretical framework of the G1–G2 scheme, including the time-local embedding scheme. In Sec. 3 we apply the derived equations to ion stopping in graphene and MoS2 monolayers using Hubbard-type lattice models in combination with a model for the resonant charge transfer between the ion and the monolayer. In Sec. 4 we present the numerical results which include the charge density dynamics in the monolayer during the ion impact. We then compute the electrostatic potential that is produced by this strongly non-uniform charge density in the the monolayer which has been found to be crucial for the emission of secondary electrons [niggas_prl_22]. Finally, we present results for the spectral properties during the ion impact. The results are summarized in Sec. 5 where we also give an outlook for future developments.
@@ 2.2 Non-equilibrium Green functions
In this section, we give a brief introduction into NEGFs and outline the derivation of the G1–G2 scheme. More information and details on the formalism can be found in Refs. [joost_prb_20, joost_prb_22, bonitz_pssb23].
@@
In this expression $h^{\mathrm{0}}(t)$ denotes the one-particle Hamiltonian, containing the kinetic energy as well as external potentials, such as the Coulomb potential of the approaching, highly-charged ion [balzer_prb16, balzer_prl_18]. $w_{ijkl}(t)$ are the matrix elements of the pair interaction potential between electrons. The time dependence reflects the adiabatic switch-on of the potential what allows us to start the simulation from a non-interacting initial state which is known [schluenzen_jpcm_19]. The definition of the one-particle Green function is given by
@@
The equations of motion for $G$ on the contour are the Keldysh-Kadanoff-Baym equations (KBEs) [keldysh64, kadanoff_quantum_2018]
@@
the equation of motion of which couples to the three-particle Green function (Martin-Schwinger-hierarchy [martin_theory_1959]). The other way [Eq. (2)] eliminates the two-particle quantity in favor of the one-particle self-energy $\Sigma$. For practical applications we switch to real times and consider the correlation functions
@@
In addition, we have access to spectral properties which require the Fourier transform, with respect to the difference time $\tau=t^{\prime}-t$, of the difference, $G^{\mathrm{>}}_{ij}(t,t^{\prime})-G^{\mathrm{<}}_{ij}(t,t^{\prime})$ [stefanucci_many-body_2013, bonitz_qkt]. However, there is also a cost to this advantage, namely the computationally expensive contour-time integral on the r.h.s. of the KBEs. The computation of this integral results in a cubic scaling of the CPU time with the propagation duration, making long simulations impossible, except for very small systems.
@@ 2.3 HF-GKBA and G1–G2 scheme
$i\hbar\frac{\partial}{\partial t}G_{ij}^{\mathrm{<}}(t)-\big[\bm{h}^{\mathrm{HF}},\bm{G}^{\mathrm{<}}\big]_{ij}(t)=i\hbar\sum_{k}\int_{t_{0}}^{\mathrm{t}}\mathrm{d}\,\bar{t}\,\left[\Sigma^{\mathrm{SOA,>}}_{ik}(t,\bar{t}\,)G_{kj}^{\mathrm{<}}(\bar{t},t)-\Sigma^{\mathrm{SOA,<}}_{ik}(t,\bar{t}\,)G_{jj}^{\mathrm{>}}(\bar{t},t)\right]\,$, (5)
where we denoted matrices in orbital space by bold letters and used the short notation $\textbf{A},\textbf{B}]_{ij}:=\sum_{k}(A_{ik}B_{kj}-B_{ik}A_{kj})$. Since the evolution on the time diagonal still depends on the off-diagonal elements, under the time integral, the equation of motion (EOM) is not fully time-diagonal and the costly memory integral still occurs.
For this reason, a common way to approximate the full two-time solution is to use a reconstruction of the off-diagonal part of the one-particle Green function [lipavsky_generalized_1986]
@@
The reconstruction already represents an approximation since additional terms, involving time integrals as well as the self-energy, are neglected [lipavsky_generalized_1986]. The ansatz (6) for the off-diagonal parts is commonly known as the Generalized Kadanoff–Baym Ansatz (GKBA). To apply this reconstruction, we also need to know the time-off-diagonal elements of the propagators $G^{\mathrm{R/A}}(t,t^{\prime})$. The most common choices are either the free GKBA or the HF-GKBA, where the propagators are the free and the Hartree-Fock propagators, respectively. In SOA, this approximation reduces the computational cost significantly (to quadratic in the number of time steps), but for higher-order self-energies the cubic scaling remains. Furthermore, due to the use of HF propagators, spectral properties do not contain correlation effects.
As described in Ref. [schluenzen_prl_20] and further expanded upon in Refs. [bonitz_pssb23, joost_prb_20, joost_prb_22], one can, in fact, by using the GKBA, completely remove the memory integral from the EOM for the Green function on the time diagonal. This is achieved by utilizing the KBEs with collision integrals defined in terms of the two-particle Green function, Eq. (1), instead of the self-energy formulation of Eq. (2). The strategy is fairly straightforward:
@@
The result is the G1–G2 scheme which has been proven to exhibit linear scaling in the propagation time at the cost of propagating, in addition to $G^{<}$, the two-particle Green function [schluenzen_prl_20].
@@
where, in Eq. (9), we introduced the two-particle Hartree and Fock Green functions. Finally, by applying the HF-GKBA and differentiating the expression of $G^{\mathrm{(2)}}$ with respect to time we obtain the coupled set of EOMs for the G1–G2 scheme in SOA [joost_phd_2022]
@@
Using the same strategy, one can find EOMs for all higher-order self-energies, as demonstrated in [joost_prb_22]. There, comparisons between the RDM and the G1–G2 approach were performed. The authors conclude that both methods are equivalent in HF, SOA and TPP approximation with and without exchange, meaning that for all known NEGF self-energies one can find an equivalent G1–G2 formulation. Since approximations stemming from self-energies can be better understood in terms of the diagrams they include, it is very useful to be able to assign them to the time-diagonal contributions. At the same time, the G1–G2 scheme allows for additional many-body approximations for which no direct single-particle self-energies are known. An example is the dynamically screened ladder (DSL) approximation, for details, see Ref. [joost_prb_22].
In this work we will focus on the HF and SOA self-energies and make use of the G1–G2 framework which has been tested multiple times for simple Hubbard chains and more complex finite graphene clusters [joost_prb_20, joost_prb_22].
@@ 2.4 NEGF embedding approach
In the following, we focus on a special case of the G1–G2 scheme, namely the description of a total system consisting of two distinct parts which will be called “system” and “environment”. Possible applications for such a framework are the description of electron dynamics in a crystal with metal contacts or, as we will discuss later in Sec. 3, an ion colliding with a graphene or MoS2 monolayer. Of course the G1–G2 scheme could be used to describe the full system in detail but that might require unnecessary computation cost since the system and the environment might evolve on completely different time scales and correlation effects might not be of equal importance in the two parts. The goal of the embedding scheme is, therefore, to be able to employ different approximations for the correlations in both, the system and the environment. For example, in the ion stopping scenario, one might want to use a higher-order self-energy, such as SOA, in the system (the target), while treating the ion without correlations altogether. The derivation of such a framework from the NEGF perspective is demonstrated, e.g., in Refs. [stefanucci_many-body_2013, balzer_cpp_21, balzer_prb_23].